10x + 3y = 8
2x + y = 2

Answer by JellyBeans(8)

10x + 3y = 8 will be Equation 1
2x + y = 2 will be Equation 2
Equation (2) 2x+y=2
y=-2x+2
Substitute the expression into equation (1)
Equation (1): 10x+3(-2x+2)=8
10x-6x+6=8
4x+6=8
4x=2
x=1/2
Substitute the answer to the expression obtained in equation (2)
y=-2x+2
y=-2(1/2)+2
y= 1
To check:
Eq (1): 10x + 3y = 8
10(1/2)+3(1)=8
5+3=8
8=8
Eq (2): 2x + y = 2
2(1/2)+1=2
1+1=2
2=2

Btw, I'm 12 years old.